# Carrer Pic de Peguera 13, 17003 Girona, Spain

### v1

@@ Abstract
The combined use of event-based vision and Spiking Neural Networks (SNNs) is expected to significantly impact robotics, particularly in tasks like visual odometry and obstacle avoidance. While existing real-world event-based datasets for optical flow prediction, typically captured with Unmanned Aerial Vehicles (UAVs), offer valuable insights, they are limited in diversity, scalability, and are challenging to collect. Moreover, there is a notable lack of labelled datasets for underwater applications, which hinders the integration of event-based vision with Autonomous Underwater Vehicles (AUVs). To address this, synthetic datasets could provide a scalable solution while bridging the gap between simulation and reality. In this work, we introduce eStonefish-scenes, a synthetic event-based optical flow dataset based on the Stonefish simulator. Along with the dataset, we present a data generation pipeline that enables the creation of customizable underwater environments. This pipeline allows for simulating dynamic scenarios, such as biologically inspired schools of fish exhibiting realistic motion patterns, including obstacle avoidance and reactive navigation around corals. Additionally, we introduce a scene generator that can build realistic reef seabeds by randomly distributing coral across the terrain. To streamline data accessibility, we present eWiz, a comprehensive library designed for processing event-based data, offering tools for data loading, augmentation, visualization, encoding, and training data generation, along with loss functions and performance metrics.
@@
spiking neural networks, event-based sensor, synthetic datasets, optical flow, underwater datasets, event-based cameras, events augmentations
@@ 1 Introduction
Event-based Cameras (EBCs), also known as dynamic vision sensors or neuromorphic cameras, are imaging sensors that operate differently from traditional frame-based cameras. Inspired by the exceptional motion perception abilities of winged insects, these cameras respond asynchronously to brightness or intensity changes with microsecond resolution (Lichtsteiner et al., 2008; Brandli et al., 2014). EBCs offer significant advantages over traditional frame-based cameras: no motion blur, high dynamic range, high temporal resolution, and low latency (Gallego et al., 2020). These unique properties make the EBC an ideal sensor to analyse dynamic scenes characterized by fast motion and rapid changes in lighting conditions.
@@
EBCs and optical flow prediction are two complementary concepts that play an integral part in motion analysis in computer vision. Traditionally, optical flow prediction has relied on analysing a sequence of frames to infer motion patterns (Shah and Xuezhi, 2021). However, EBCs present a new paradigm by providing an asynchronous and sparse representation of data in the form of a stream of events (Gallego et al., 2020). As such, many methodologies have been developed for event-based processing, though primarily for relatively simple scenes containing well-defined shapes, edges, and corners (Kueng et al., 2016; Zhu et al., 2017; Gehrig et al., 2020). However, these methods have proven to be inefficient and computationally expensive in complex scenes.
@@
An alternative approach is the use of Neural Networks (NNs) for processing event-based data, which has demonstrated remarkable success in modelling and understanding the nature of these biological vision systems. One significant drawback of EBCs, however, is their incompatibility with conventional NN models, such as Convolutional Neural Networks (CNNs). This requires the use of complex data encoding schemes to convert the sparse asynchronous stream of events into a frame-based representation that can be handled by CNNs, as demonstrated by researchers (Zhu et al., 2018b, 2019; Wan et al., 2022) who have implemented U-net (Weng and Zhu, 2015) like architectures for event-based optical flow estimation. However, such compression of event streams into frame-based representations results in loss of information, thereby limiting the potential of EBCs. As a result, there is a demand for newer paradigms specifically optimized for handling such data. An SNN is one such brain-inspired computing model that relies on so-called spiking neurons to process information (Tavanaei et al., 2019). Spiking neurons are inherently designed to process temporal patterns, making them naturally compatible with the asynchronous stream of events. An interconnection of such spiking neurons in an SNN allows them to effectively extract spatial features, enabling a comprehensive framework for encoding dynamic visual motion over time.
@@
Nonetheless, there is still a lack of sufficiently diverse and well-labelled datasets to facilitate model training. Generating real-world data is both expensive and time-consuming, often requiring specialized equipment to obtain corresponding ground truth data as an alternative to the tedious process of manual annotations by human experts. For instance, datasets such as MVSEC (Zhu et al., 2018a) and DSEC (Gehrig et al., 2021), utilize 3D LiDARs to create a 3D map of the environment. This reconstructed environment is then projected onto the EBC sensor to generate a dense optical flow for use as ground truth during training. However, the resulting optical flow includes some inaccuracies due to clock synchronization issues and often tends to miss information, failing to fully represent the event sequences captured by the sensor. As such, these real-world datasets may not be ideal for model training or evaluation, which stresses the need for a synthetic alternative. In fact, due to the nature of EBCs, synthetic event-based data may translate well to real-world data. The features extracted from synthetic frame-based camera datasets often exhibit obvious differences from the features extracted from real-world data. These differences typically arise from distinct lighting conditions, camera noise, and environmental complexity in real-world settings which are challenging to replicate accurately in synthetic data. Such variations can lead to shifts in pixel intensities, colour distribution and texture details, resulting in significant differences between the features extracted from synthetic and real-world images. In contrast, EBCs only trigger events when detecting brightness changes, producing an image of events that only has 2 values: a positive or a negative polarity. As such, unlike frame-based images that rely on a static grid of pixels with values typically ranging between 0 and 255, the output of EBCs is simpler, leaving more room for error when transitioning from simulation.
@@
To our knowledge, no event-based optical flow prediction datasets currently exist for underwater environments. Capturing such data in real-world conditions presents significant challenges due to several factors such as poor lighting conditions, water turbidity, and light scattering. These factors adversely affect the performance of conventional camera systems typically used for capturing optical flow data. To overcome these limitations, we take a synthetic approach, enabling precise control over environmental conditions and ensuring reliable ground truth. A synthetic event-based dataset for underwater optical flow prediction opens new avenues for research on EBCs in these conditions. Light absorption and scattering, especially at greater depths, reduce illumination and cause colour distortion. Unlike conventional cameras, EBCs operate based on changes rather than ambient light, allowing them to function effectively in low-light scenarios. Moreover, EBCs are highly resilient to motion blur and can effectively filter out noise caused by scattering effects, making them ideal for capturing fast-moving objects in dynamic underwater environments. Their high dynamic range further enables them to overcome the contrast limitations of conventional frame-based cameras.
@@
As such, we present eStonefish-scenes, a synthetic event-based optical flow dataset tailored specifically for underwater environments, together with an easy-to-use pipeline for creating custom datasets. eStonefish-scenes is based on the Stonefish simulator (Grimaldi et al., 2025), with a Remotely Operated Vehicle (ROV) capturing event-based, grayscale, and optical flow data. To further enhance the realism of our simulations, we introduce Stonefish-scenegen, a scene generator capable of randomly populating the environment with corals, creating rich and vibrant life-like reefs. To animate these scenes, we also present Stonefish-boids, a package which simulates the behaviour of schools of fish, adding realistic dynamics to the environment. Together, these components enable the creation of highly realistic underwater scenarios. The dataset is hosted on Zenodo (10.5281/zenodo.15130453) and the data generation pipeline is publicly available at the following link: https://github.com/CIRS-Girona/estonefish-scenes.
@@
Finally, we also introduce eWiz, a library that is composed of all the tools necessary for processing event-based data, ranging from data loading and manipulation, to augmentation and visualization, to encoding and training data generation, together with various loss functions and performance metrics. We plan to add more functionalities and incorporate more areas of applications as our work on event-based processing progresses. The library is openly available at the following link: https://github.com/CIRS-Girona/ewiz.
In summary, our main contributions are:
@@
  • The creation of eStonefish-scenes, an event-based optical flow dataset specifically tailored for underwater environments, along with an open data generation pipeline, which simplifies the process of extending the datasets. To the best of our knowledge, no prior event-based datasets exist for underwater applications.
@@
The remainder of this paper is organized as follows. Section 2 presents an overview of other publicly available datasets for event-based optical flow prediction. Section 3 and Section 4 describe the details of the simulation setup and sensor configurations used for the dataset, along with the data generation pipeline. Section 5, on the other hand, presents details on the eWiz library for additional processing of the datasets for visualization, analysis and model training. Section 6 gives an overview of how the datasets are structured. Finally, Section 7 concludes this study and presents directions for future work.
@@ 2 Related Work
In the literature, only a limited number of real-world event-based datasets have been employed for developing and evaluating event-vision algorithms and NNs targeting optical flow prediction. Among these, the widely used MVSEC dataset (Zhu et al., 2018a) stands out, offering two distinct scenarios: a stereo pair of EBCs mounted on an UAV in an indoor environment, and the same setup mounted on a car in an outdoor setting. The ground truth optical flow for indoor scenes is derived using a motion capture system, while for outdoor scenes, it is obtained through a combination of GPS data and scene reconstruction techniques. Building upon MVSEC, the DSEC dataset, introduced by Gehrig et al. (2021), presents significant improvements. It features driving scenarios with displacement field magnitudes reaching up to 210 pixels and a camera resolution three times higher than that of MVSEC. While these datasets have been instrumental in advancing event-based vision research, they are not without limitations. The generation of real-world data is both resource-intensive and time-consuming, often necessitating specialized equipment to capture accurate ground truth information. This reliance on costly setups further underscores the challenges of producing high-quality datasets in this domain. Moreover, the MVSEC dataset exhibits notable limitations in the accuracy of its ground truth optical flow. The ground truth was generated using a Velodyne LiDAR, which relies on a 3D reconstruction of the environment to estimate optical flow. The dataset suffers from additional challenges due to the mechanical setup. Specifically, the vibrations of the car during forward motion induce a wobbling effect in the event-based sensor. This wobble results in the sensor capturing significant up-and-down motion, whereas the ground truth optical flow predominantly reflects forward motion. This mismatch between the motion perceived by the event-based sensor and the ground truth optical flow can introduce discrepancies in training and evaluating models, especially for applications that rely on precise motion estimation.
@@
These challenges contribute to the scarcity of real-world event-based datasets, prompting researchers to increasingly turn to synthetically generated datasets. Leveraging advanced simulators, synthetic datasets provide a scalable and cost-effective alternative, enabling the generation of diverse and accurately labelled data. ESIM (Rebecq et al., 2018) for example, being one of the most accurate in the literature, makes use of an adaptive rendering scheme that only samples frames when necessary. Such a rendering scheme adapts the sampling rate to the predicted dynamics of the visual signal, allowing for a faithful simulation of events when fast motion occurs. ESIM has been used to generate several synthetic event-based optical flow datasets, including MultiFlow (Gehrig et al., 2024), and HREM (Luo et al., 2024). MultiFlow contains multiple moving objects undergoing continuous similarity transformations, along with dense ground-truth optical flow. HREM, on the other hand, comprises 100 virtual scenes designed to replicate real-world indoor and outdoor environments. BlinkSim, proposed by Li et al. (2023), is another EBC simulator that includes a powerful rendering engine capable of building thousands of scenes with different objects and motion patterns at high frequency. This simulator gave rise to the BlinkFlow dataset, which improves generalization performance of state-of-the-art methods by 40% on average. Another synthetic dataset is the MDR created by Luo et al. (2023). It makes use of Blender to create indoor and outdoor 3D scenes and includes diverse camera motions and high frame-rate videos between images.
@@
One of the reasons behind generating synthetic datasets is the ability to diversify the created environments and easily extend applications related to EBCs. Although synthetic optical flow datasets for event-based vision are already available, most are not specifically designed to reflect the motion dynamics or characteristics of any particular vehicle type. MDR, ESIM, and BlinkFlow are limited to predefined camera trajectories, which restricts their ability to accurately replicate such dynamic motion patterns. Furthermore, these simulators lack sophisticated light behaviour modelling, which is essential for creating realistic and complex environments. For example, factors such as absorption, scattering, and refractive effects, key characteristics of underwater environments, that these simulators fail to account for, undermine the fidelity of synthetic datasets for underwater applications. On the other hand, ESIM makes use of an OpenGL rendering engine (The Khronos Group, Inc., 2014). Similarly, the use of Blender The Blender Foundation (2002) for MDR requires extensive expertise in 3D scene creation to appropriately simulate real-world lighting conditions, especially for underwater environments. All these factors, make the creation of synthetic event-based datasets a challenging endeavour, particularly for underwater environments.
@@
Currently, research on underwater event-based applications is highly limited, especially due to the lack of publicly available datasets captured with EBCs in underwater environments. To address this gap, we created a diverse synthetic dataset tailored for underwater applications, along with a modular data generation pipeline to facilitate creation of custom datasets. The dataset incorporates both static and dynamic elements, enabling the trained algorithms to learn robust feature representations by capturing background stability and motion-driven variations. This would also enhance generalization to real-world scenarios, where stationary structures and moving entities, such as marine life, coexist.
@@
Finally, although a few open-source tools are available for specific event-based data processing tasks, such as the work by Shiba et al. (2022), no comprehensive library currently integrates these capabilities into a unified framework. To address this gap and streamline future development, we developed eWiz, a comprehensive library for event-based data manipulation and processing.

### v2

@@ Abstract
Event-based cameras (EBCs) are poised to transform underwater robotics, yet the absence of labelled event-based datasets for underwater environments severely limits progress in tasks such as visual odometry and obstacle avoidance. Real-world event-based optical flow datasets are scarce, resource-intensive to collect, and lack diversity, while no prior benchmarks target underwater applications. To bridge this gap, we introduce eStonefish-Scenes, a synthetic event-based optical flow dataset generated using the Stonefish simulator, together with an open data generation pipeline for creating customizable underwater environments featuring realistic coral reefs and biologically inspired schools of fish with reactive navigation behaviours. We also present eWiz, a comprehensive library for event-based data processing, encompassing data loading, augmentation, visualization, encoding, training utilities, loss functions, and evaluation metrics.
To validate sim-to-real transferability, we collected real-world data using a DAVIS346 hybrid event-and-frame camera mounted on a BlueROV2 in an indoor testing pool. Ground-truth optical flow was derived via homography-based frame-to-poster registration, and per-pixel uncertainty was estimated through Monte Carlo perturbation of keypoint correspondences. This uncertainty was incorporated into the evaluation metrics, enabling reliability-aware performance assessment. A ConvGRU-based optical flow network, trained exclusively on synthetic eStonefish-Scenes data, was evaluated on the real-world sequences without fine-tuning, achieving an uncertainty-weighted average endpoint error of 0.79 pixels. These results demonstrate that the proposed synthetic dataset effectively supports sim-to-real transfer for underwater event-based optical flow estimation, substantially reducing the need for costly real-world data collection.
@@
event-based sensor, synthetic datasets, optical flow, underwater datasets, event-based cameras, events augmentations, neural networks, data validation, sim-to-real transfer performance, event-based datasets
@@ 1 Introduction
Event-based Cameras (EBCs), also known as dynamic vision sensors or neuromorphic cameras, are imaging sensors that operate differently from traditional frame-based cameras. Inspired by the motion perception of winged insects, these cameras respond asynchronously to brightness changes with microsecond temporal resolution (camera1Lichtsteiner08; camera2Brandli14). EBCs offer significant advantages over traditional frame-based cameras: virtually no motion blur, high dynamic range, high temporal resolution, and low latency (camera3Gallego20). These unique properties make the EBC an ideal sensor to analyse dynamic scenes characterized by fast motion and rapid changes in lighting conditions.
@@
EBCs and optical flow prediction are two complementary concepts that play an integral part in motion analysis in computer vision. Traditionally, optical flow prediction has relied on analysing a sequence of frames to infer motion patterns (flowShah21). However, EBCs present a new paradigm by providing an asynchronous and sparse representation of data in the form of a stream of events (camera3Gallego20). As such, many methodologies have been developed for event-based processing, aimed primarily at relatively simple scenes containing well-defined shapes, edges, and corners (icpKueng16; probabilityZhu17; ekltGehrig20). However, these methods have proven to be inefficient and computationally expensive in complex scenes (trackingMessikommer23).
@@
An alternative approach is the use of Neural Networks (NNs) for processing event-based data, which has demonstrated remarkable success in modelling and understanding the nature of biological vision systems. One significant drawback of EBCs, however, is their incompatibility with conventional NN models, such as Convolutional Neural Networks (CNNs). This requires complex data encoding schemes to convert the sparse asynchronous stream of events into a frame-based representation that can be handled by CNNs. This approach has been demonstrated by researchers (evflownetZhu18; egoflownetZihao19; biflowWan22) who implemented U-Net-like architectures (unetRonneberger15; unetWeng15) for event-based optical flow estimation. Compression of event streams into frame-based representations results in loss of information, thereby limiting the full potential of EBCs. As a result, there is a demand for newer paradigms specifically optimized for handling such data. An Spiking Neural Network (SNN) is one such brain-inspired computing model that relies on so-called spiking neurons to process information (snnTavanaei19). Spiking neurons are inherently designed to process temporal patterns, making them naturally compatible with the asynchronous stream of events. The interconnection of spiking neurons in an SNN allows them to effectively extract spatial features and encode dynamic visual motion over time.
@@
Sufficiently diverse and well-labelled datasets to facilitate model training remain scarce. Generating real-world data is both expensive and time-consuming, often requiring specialized equipment to obtain corresponding ground truth data as an alternative to the tedious process of manual annotations by human experts. For instance, datasets such as MVSEC (mvsecZhu18) and DSEC (dsecGehrig21) utilize 3D LiDARs to create a 3D map of the environment. The reconstructed environment is then projected onto the EBC sensor to generate a dense optical flow for use as ground truth during training. However, the resulting optical flow is affected by clock synchronization issues and incomplete coverage, failing to fully represent the event sequences captured by the sensor. As such, these real-world datasets may not be ideal for model training or evaluation, which stresses the need for a synthetic alternative.
Recently, the landscape of underwater event-based datasets has expanded with synthetic benchmarks such as UEOF (ueofTruong26) and AquaticVision (aquaticvisionPeng25). UEOF provides a high-resolution benchmark derived from physically-based ray-traced RGBD sequences, utilizing a video-to-event (v2e) conversion pipeline to generate pseudo-events. However, this approach has notable limitations: it relies on post-processing pre-rendered video frames rather than natively integrating the event sensor within the simulator, and it requires manually crafted camera trajectories that are decoupled from any underlying robot dynamics. This decoupling prevents the dataset from capturing the authentic interplay between vehicle motion and sensor response characteristic of real underwater robotics scenarios. In contrast, our approach integrates the event camera directly into the Stonefish simulator, ensuring that the generated data captures the authentic coupling between the robot’s hydrodynamics (e.g., heave, surge, sway) and the sensor’s temporal response. Furthermore, unlike datasets that target high-resolution benchmarks, eStonefish-Scenes deliberately retains the native resolution of the DAVIS346 sensor ($346\times 260$ pixels). This design choice is critical for enabling the deployment of lightweight, resource-efficient neural networks capable of running on the power-constrained embedded hardware typical of AUVs. Finally, whereas existing benchmarks often lack a direct sim-to-real validation component, we validate our synthetic data against our own dedicated real-world dataset collected with a similar sensor configuration.
In fact, due to the nature of EBCs, synthetic event-based data can translate well to real-world data. The features extracted from synthetic frame-based camera datasets often exhibit obvious differences from the features extracted from real-world data. These differences typically arise from distinct lighting conditions, camera noise, and environmental complexity in real-world settings, which are challenging to replicate accurately in synthetic data. Such variations can lead to shifts in pixel intensities, colour distribution and texture details, resulting in significant differences between the features extracted from synthetic and real-world images. In contrast, EBCs only trigger events when detecting brightness changes, producing an event representation that only has two values: a positive or a negative polarity. As such, unlike frame-based images that rely on a static grid of pixels with values typically ranging between 0 and 255, the output of EBCs is simpler, allowing for greater tolerance to domain differences when transitioning from simulation.
@@
To the best of our knowledge, no event-based optical flow prediction datasets currently exist for underwater environments. Capturing such data in real-world conditions presents significant challenges due to several factors, such as poor lighting conditions, water turbidity, and light scattering. These factors adversely affect the performance of conventional camera systems typically used for capturing optical flow data. To overcome these limitations, we take a synthetic approach, enabling precise control over environmental conditions and ensuring reliable ground truth. A synthetic event-based dataset for underwater optical flow prediction opens new avenues for research on EBCs in these conditions. Light absorption and scattering, especially at greater depths, reduce illumination and cause colour distortion. Unlike conventional cameras, EBCs respond to changes rather than intensities of ambient light, allowing them to function effectively in low-light scenarios. Moreover, EBCs are resilient to motion blur and noise caused by scattering effects, making them ideal for capturing fast-moving objects in dynamic underwater environments.
@@
To address this need, we present eStonefish-Scenes, a synthetic event-based optical flow dataset tailored specifically for underwater environments, together with an easy-to-use pipeline for creating custom datasets. eStonefish-Scenes is based on the Stonefish simulator (stonefishCielak19; stonefishGrimaldi25), with a Remotely Operated Vehicle (ROV) capturing event-based, grayscale, and optical flow data. To enhance the realism of the simulations, we introduce Stonefish-SceneGen, a scene generator capable of randomly populating the environment with corals, creating rich and vibrant lifelike reefs. To animate these scenes, we also present Stonefish-Boids, a package that simulates the behaviour of schools of fish. Together, these components enable the creation of realistic underwater scenarios. The dataset is hosted on Zenodo (https://doi.org/10.5281/zenodo.15130452) and the data generation pipeline is publicly available at: https://github.com/CIRS-Girona/stonefish-scenegen.
@@
We also introduce eWiz, a library comprising of all the tools necessary for processing event-based data, ranging from data loading and manipulation, to augmentation and visualization, to encoding and training data generation, together with various loss functions and performance metrics. The library is openly available at the following link: https://github.com/CIRS-Girona/ewiz.
To evaluate the sim-to-real transfer performance of the synthetic dataset, a real-world experimental study was conducted. A DAVIS346 event-based camera was rigidly mounted on a BlueROV2 platform and deployed in the indoor testing pool at the Centre for Research in Underwater Robotics and Vision (CIRS) in Girona. To enable controlled and repeatable optical flow estimation, a large, high-resolution textured poster was placed on the pool floor, providing a planar reference surface with rich visual features. Ground-truth optical flow was computed via frame-to-poster feature matching using a homography-based registration approach (homoHartley04). To quantify the reliability of the resulting flow, per-pixel uncertainty was estimated via Monte Carlo perturbations of keypoint correspondences, yielding covariance-based uncertainty maps aligned with the optical flow field. This uncertainty information was subsequently incorporated into the evaluation metrics, enabling confidence-weighted performance assessment rather than relying on fixed outlier thresholds commonly used in optical flow benchmarks. The real-world validation dataset for eStonefish-Scenes can be found on Zenodo (https://doi.org/10.5281/zenodo.18473013).
Finally, sim-to-real generalisation was assessed by implementing a ConvGRU-based event-based optical flow network inspired by (firenetHagenaars21). Training was performed exclusively on the synthetic eStonefish-Scenes dataset and evaluation on the real-world sequences. The network was trained using self-supervised photometric and smoothness losses defined over static scenes (evflownetZhu18), and its performance on real data was measured using uncertainty-aware error metrics. This evaluation demonstrates the effectiveness of the proposed synthetic dataset for learning transferable motion representations and highlights its suitability for underwater event-based vision tasks.
The main contributions of the work are:
@@
  • The creation of eStonefish-Scenes, an event-based optical flow dataset specifically tailored for underwater environments, along with an open data generation pipeline that simplifies the process of extending the dataset. To the best of our knowledge, no prior event-based datasets exist for underwater applications.
@@
  • A real-world underwater data acquisition approach using an event-based camera mounted on an ROV, along with a homography-based method for ground-truth optical flow computation and per-pixel uncertainty estimation.
  • An uncertainty-aware evaluation framework for event-based optical flow, incorporating covariance-based confidence measures into standard performance metrics.
  • A sim-to-real validation study demonstrating the effectiveness of the proposed dataset by training a ConvGRU-based event-based optical flow network exclusively on synthetic data and evaluating it on real underwater sequences.
The remainder of this paper is organized as follows. Section 2 reviews publicly available event-based datasets for optical flow prediction. Section 3 describes the simulation setup and sensor configurations used to generate eStonefish-Scenes, while Section 4 details the data generation and collection pipeline. Section 5 presents the eWiz library for event-based processing, visualization, encoding, and training utilities. Section 6 describes the dataset structure and storage format. Section 7 introduces the real-world validation dataset acquisition methodology and the homography-based ground-truth optical flow computation with uncertainty estimation. Section 8 evaluates sim-to-real transfer by training an event-based optical flow network on synthetic data and testing on real sequences. Finally, Section 9 concludes the paper and outlines future work.
@@ 2 Related Work
In the literature, only a limited number of real-world event-based datasets have been employed for developing and evaluating event-based vision algorithms and NNs targeting optical flow prediction. Among these, the widely used MVSEC dataset (mvsecZhu18) stands out, offering two distinct scenarios: a stereo pair of EBCs mounted on a Unmanned Aerial Vehicle (UAV) in an indoor environment, and the same setup mounted on a car in an outdoor setting. The ground truth optical flow for indoor scenes is derived using a motion capture system, while for outdoor scenes, it is obtained through a combination of GPS and scene reconstruction techniques. Building upon MVSEC, the DSEC dataset, introduced by dsecGehrig21, presents significant improvements, including higher camera resolution and larger displacement fields. While these datasets have been instrumental in advancing event-based vision research, they are not without limitations. The generation of real-world data is both resource-intensive and time-consuming, often requiring specialized equipment to capture accurate ground-truth information. The MVSEC dataset exhibits notable limitations in the accuracy of its ground-truth optical flow. The ground truth was generated using a Velodyne LiDAR, which relies on a 3D reconstruction of the environment to estimate optical flow. The dataset suffers from additional challenges due to the mechanical setup. Specifically, the vibrations of the car during forward motion induce a wobbling effect in the event-based sensor. This wobbling motion leads to the sensor capturing significant vertical motion, whereas the ground truth optical flow predominantly reflects forward motion. This mismatch between the motion perceived by the event-based sensor and the ground-truth optical flow can introduce discrepancies when training and evaluating models, especially for applications that rely on precise motion estimation.
@@
These challenges contribute to the scarcity of real-world event-based datasets, prompting researchers to increasingly turn to synthetically generated datasets, produced by advanced simulators. ESIM (esimRebecq18), for example, being one of the most accurate in the literature, makes use of an adaptive rendering scheme that only samples frames when necessary. Such a rendering scheme adapts the sampling rate to the predicted dynamics of the visual signal, allowing for a faithful simulation of events when fast motion occurs. ESIM has been used to generate several synthetic event-based optical flow datasets, including MultiFlow (multiflowGehrig24) and HREM (hremLuo24). MultiFlow contains multiple moving objects undergoing continuous similarity transformations, along with dense ground-truth optical flow. HREM, on the other hand, comprises 100 virtual scenes designed to replicate real-world indoor and outdoor environments. BlinkSim, proposed by blinksimLi23, is another EBC simulator that includes a powerful rendering engine capable of generating thousands of scenes with different objects and motion patterns at high frequency. This simulator gave rise to the BlinkFlow dataset, which improves the generalization performance of state-of-the-art methods by 40% on average. Another synthetic dataset is MDR created by mdrLuo23. It makes use of Blender (blenderBlender23) to create indoor and outdoor 3D scenes and includes diverse camera motions and high-frame-rate videos.
@@
One of the key reasons behind generating synthetic datasets is the ability to diversify the created environments and easily extend applications related to EBCs. Although synthetic optical flow datasets for event-based vision are already available, most are not specifically designed to reflect the motion dynamics or characteristics of any particular vehicle type. MDR, MultiFlow, and BlinkFlow are limited to predefined camera trajectories, which restricts their ability to accurately replicate such dynamic motion patterns. Furthermore, these simulators lack sophisticated light behaviour modelling, which is essential for creating realistic and complex environments. Phenomena such as light absorption, scattering, and refracted sunlight are key characteristics of underwater environments that these simulators fail to account for, thus limiting the fidelity of synthetic datasets for underwater applications. On the other hand, ESIM makes use of an OpenGL rendering engine (openglOpengl14), which requires manual development of every rendering pipeline and shader from scratch, making it even more complex for underwater environments. Similarly, the use of Blender (blenderBlender23) for MDR requires some knowledge in 3D scene creation to appropriately simulate realistic conditions, especially for underwater environments. All these factors make the creation of synthetic event-based datasets a challenging endeavour, particularly for underwater environments.
@@
Currently, research on underwater event-based applications is highly limited, especially due to the lack of publicly available datasets captured with EBCs in underwater environments. To address this gap, we created a diverse synthetic dataset tailored for underwater applications, along with a modular data generation pipeline to facilitate the creation of custom datasets. The dataset incorporates both static and dynamic elements, enabling the trained algorithms to learn robust feature representations by capturing background stability and motion-driven variations. This also enhances generalization to real-world scenarios, where stationary structures and moving entities, such as marine life, coexist.
@@
Finally, although a few open-source tools are available for specific event-based data processing tasks, such as the work by mcShiba2022, no comprehensive library currently integrates these capabilities into a unified framework. To address this gap and streamline future development, we developed eWiz, a comprehensive library for event-based data manipulation and processing.